OpenStudy (anonymous):

Which of the following sets of numbers is not a Pythagorean triplet? 12, 16, 20 14, 48, 50 16, 30, 34 20, 48, 56

OpenStudy (anonymous):

You can check each by simply plugging the values into the pythagorean theorem : a^2+b^2=c^2 In this case, the last set of numbers is not the pythagorean triplet. It comes out to 20^2+48^2=?= 6^2 2704=/=3136 which is not true

OpenStudy (anonymous):

thanks